2. Let A be an n x n matrix with AT =-A (a) Prove that A has value 0. (b) Prove that A has determinant 0 if n is odd.
Answer #1

0 let Abe an aon matrin with AT =-A - A is skew symmetric ATS-A & A is sxew symmetric aija-aji & ith yow and jth column. (a)a matrix is the product of air the value of but air=0 therefore, value of A=0 (6) to 111=0 is odd. we need prove that it an sdet(A) = det (AT) = det(A) (since 17 =-^) = (-1) det (A) della) = det (a).(-1) -det(a) (sina, nisool) det(A) = - det (A) -)